OpenStudy (anonymous):

how do You Calculate the volume of a sphere where the radius is 9 meters?

OpenStudy (anonymous):

Volume of a sphere \(= \frac 43 \times \pi \times r^3 \)

OpenStudy (turingtest):

the formula for the volume of a sphere is \[\large V=\frac43\pi r^3\]lol I thought you'd give some crazy formula FFM

OpenStudy (turingtest):

I bet you know 100 ways to find the volume of a shpere

OpenStudy (anonymous):

Hehe, not 100 (definitely) ;)

OpenStudy (anonymous):

Thank You !

OpenStudy (turingtest):

welcome!